-4x - x - 3(x - 2) = 4 - (x - 2)
Answer by checkley77(12844)

-4x - x - 3(x - 2) = 4 - (x - 2)
-4X-X-3X+6=4-X+2
-8X+X=6-6
-7X=0
X=0 ANSWER.
PROOF:
-4*0-0-3(0-2)=4-(0-2)
0-0-3*-2=4+2
6=6